### v2

Abstract
Organic electro-optic materials have demonstrated promising performance in developing electro-optic phase shifters. Their integration with other silicon photonic processes, nanofabrication complexities, and durability remain to be developed. While the required poling step in electro-optic polymers limits their potential and utilization on a large scale, devices made of paraelectric nematic liquid crystals suffer from slow bandwidth. In ferroelectric nematic liquid crystals, we report an additional GHz-fast phase shift that ultimately allows for significant second-order nonlinear optical coefficients and the related Pockels effect. It avoids poling issues and can pave the way for hybrid silicon-organic systems with CMOS-foundry compatibility. We report DC and AC modulation efficiencies of 0.25 Vmm and 25.7 Vmm, respectively, an on-chip insertion loss of 2.1 dB, and an electro-optic bandwidth of >4.18 GHz, employing improved light-matter interaction in a waveguide architecture that calls for only one lithography step.
keywords:
Silicon organic hybrid material, electro optic polymer, liquid crystal, phase shifter, modulatorIntroduction
Silicon and silicon nitride photonics have enabled several recent technological advancements, and future roadmaps have been designed based on their capabilities and integration with established CMOS foundries [1]. The electro-optic phase shifter (EOPS) [2] is one of the most fundamental components upon which other essential building blocks in photonic integrated circuits (PIC) are based, including electro-optic (EO) switches [3], EO modulators [4], optical interconnects [5], and comb generators [6]. Traditional phase shift mechanisms often show poor overall performance, such as limited EO bandwidth (BW) and high static energy consumption () when leveraging the thermo-optic effect or free-carrier dispersion effect with traveling wave modulators. Free carrier dispersion also leads to doping-induced insertion loss (IL). Innovations in hybrid material platforms promise viable solutions to traditional Si photonics (SiP) shortcomings. However, they typically face significant challenges, such as several complex nanofabrication stages, scaling, and interaction with current CMOS foundries. Modulators based on indium tin oxide (ITO) have recently experienced notable improvements [4,7-9]. The additional operations necessary for ITO/insulator thin-film deposition, demanding quality control methods, and their low extinction ratio (ER) suggest further R&D needs [10]. Carrier-free phase shift processes, such as the Pockels effect, are among the more appealing techniques, especially regarding EO bandwidth and energy consumption. These metrics and low IL can excel in LiNbO3 on silicon (LNOI) [11-12]. Due to the low EO coefficient of LiNbO3, however, phase shifters based on LNOI require either a large footprint or a high drive voltage (), where is the voltage necessary to produce shift. A material system based on barium titanate (BTO) can simultaneously meet the requirements of low IL and high EO bandwidth [13-14]. Despite having a significant EO coefficient, BTO is unable to fully utilize the improved light-matter interactions (LMI) provided by some waveguide topologies, such as slot waveguides, failing to give the record-low modulation efficiencies (), where is the EOPS effective length [15]. BTO also involves more meticulous fabrication steps, loses its high EO efficiency at cryogenic temperatures, and necessitates process controls for optimum domain development [16].
Furthermore, the preceding approaches cannot meet the trade-offs between various device metrics of a particular modulator, such as footprint, EO and optical BW, IL, , , and dynamic energy consumption () [17-18]. Hybrid integration of Si with organic electro-optic (OEO) materials has been proposed to overcome these shortcomings. They have advantages over their inorganic equivalents due to a combination of their EO and physicochemical characteristics. The EO properties of OEO materials, including the Pockels factor () and optical loss, have continuously progressed over the last two decades ( as large as 9000 pmV-1 and absorption coefficients as low as have been reported [19], where is the OEO refractive index and represents the Pockels coefficient). Thanks to their liquid state, an OEO can infiltrate nanostructures during the coating step and benefit from the enhanced LMI. Along with the resultant high modulation efficiencies, the low dielectric dispersion [20] of OEO promises better modulators. As a result, <1 V, sub-mm-long Mach-Zehnder modulators (MZM) with IL < 1 dB, EO bandwidth surpassing 100 GHz, and 100 Jbit-1 have been enabled [21]. Such an all-inclusive, miniaturized modulator can work in extreme temperatures (as high as 383 K (110C) [22] and as low as 4 K [23,24]), which encompass the crucial requirements for a state-of-the-art modulator.
While EO polymers have attracted increased attention mainly because of their remarkable EO characteristics, their application to photonic-integrated circuits (PIC) has challenges. Multiple process steps and often hazardous solvents are needed to prepare EO polymer thin films, necessitating equipment for precise mixing, coating, and drying. In addition, to align chromophores in an EO polymer material, an electro-thermal poling step is necessary. On a larger scale, dedicated DC lines may be required depending on the waveguide geometry and accompanying LMI. Despite considerable breakthroughs in producing thermally and temporally stable polymers, not all poled polymers are resilient to high temperatures ([22]), aging ([25]), and poling-induced optical loss ([26]). For practical applications, the devices may still require a certain level of hermetic sealing to become temperature- and humidity-resistant. EO polymers are also recognized for their lower resistance at higher temperatures, which coincides with detrimental leakage currents. It leads to decreased EO efficiency, which has a direct impact on device performance. Additional nanofabrication steps have been identified as advantageous to resolving this issue, such as exploiting a thin charge-barrier layer comprised of alumina or titania created by atomic layer deposition, [27,28]. Poling also limits high-temperature excursions typical of back-end-of-line processing since such excursions would destroy the poling-induced order. The problems above could increase the overall cost, add complexity, and jeopardize the integration of polymer-based hybrid platforms.
Another class of OEO material is PN-LC, which provides a variety of efficient but slow phase shift processes, such as the birefringence effect [29-32]. Ultra-high modulation efficiencies have been reported using strong-LMI implementations, including infiltrated slot waveguides with the material [2]. Unlike polymers, preparing PN-LC devices does not involve a solvent that must be evaporated during post-baking processes. It simplifies the spin coating process and may allow for alternative coating approaches. The material, for example, can be coated on photonic parts via microfluidic channels or fluid dispensing. A PN-LC-based platform is attractive for realizing EO switches, phase shifters, or shutters, but it is not desirable for high-speed applications.
Another state of matter, ferroelectric nematic liquid crystals (FN-LC), has recently been discovered [33-35]. In the landscape of organic EO materials, FN-LC is situated between poled polymers, organic nonlinear optic (NLO) crystals, and PN-LCs, bringing the most desirable aspects of each (see Table 1). Like poled polymers and organic NLO crystals, FN-LC is non-centrosymmetric and has nonzero second-order susceptibilities (). Like organic NLO crystals and PN-LCs, they organize spontaneously–the ordered state is the thermodynamic minimum with significant order parameters. Similar to PN-LCs, the alignment can be controlled over large areas by applying alignment layers at the interface or small external fields. They can also be switched like PN-LCs, with a reorientation of the fast and slow axes in an external field. A unique feature of the FN-LC is that the molecular long axis (slow axis) is highly polarizable and coincident with the director and the polar axis (Fig. 1(a) inset). This creates the condition for achieving very large spontaneous polarizations and large second-order nonlinearities, which have been observed by second harmonic generation measurements [36]. These nonlinearities have been dramatically enhanced by adding NLO dyes to the polar matrix of the FN-LC host material [37]. However, what remains to be demonstrated is a fast EO effect (i.e., Pockels effect) observed in other non-centrosymmetric organic materials. Various operational regimes for different variations of FN-LC were described in [38], where the authors explored other EO properties of the material but did not report the sub-nanosecond characteristic relaxation times or the Pockels effect we demonstrated here. The study found that may be attained with mild external electric fields (). At larger values over 1 V, their model predicted a smaller relaxation time inversely proportional to , down to .
We report the proofs of a considerably faster EO effect in FN-LC, utilizing the nanoscale properties of our SOH structure. We specifically show that FN-LC can be integrated with similar ease as PN-LCs onto a SiP platform, that alignment of the polar axis can be easily achieved over the length of the device, and that a significant EO response can be observed. This paper proposes a hybrid material platform for PICs comprising numerous moderate-speed EOPS that prioritize large-scale scaling, integrability, and ease of fabrication. This includes non-modulation applications, such as switches [3,39], transducers [40], and weight banks [41] for neuromorphic PICs [42] and quantum computing [43].
Results
LMI-engineered design
Waveguides such as Si [44] and plasmonic [45] slotted, sub-wavelength grating (SWG) [46], photonic-crystal [47], and thinned [22], are traditionally employed as EOPS in modulators. However, they often add excessive loss [48], limit optical BW [48], suffer from poor infiltration, and poling [27,49]. Of particular interest for the current work is the extra fabrication stages slotted-based devices demand [40], some of which are not quite CMOS-compatible. For instance, opening oxide cladding over isolated Si islands of SWG or metallic slots of plasmonic waveguides may violate design rules in wafer nanofabrication. Also, exploiting the benefits of structures such as thinned waveguides necessitates taper converters from complete to partial etch Si [22]. Geometries required for efficient and low-loss mode converters, such as strip-to-slot, are not always accessible through nanofabrication runs or have integration challenges with the rest of the PIC. On the other hand, a non-slotted waveguide does not have these issues at the cost of offering a weak LMI [40], thus requiring a high voltage for poling and driving. The device size and EO bandwidth must then be compromised to keep the voltage within the ranges supplied by CMOS drivers.
We propose a finger-loaded non-slotted (FNS) waveguide consisting of two strip-to-finger mode converters to resolve this trade-off, as seen in Fig. 1(a). In analogy to segmented slotted [50] and non-slotted [51] topologies, this structure eliminates the misalignment problem seen in strip-loaded slot waveguides by offering at least one fewer step for lithography and etching. The periodicity and duty cycle of the subwavelength structure of the FNS waveguide are restricted by fabrication capabilities yet selected to produce a better balance between electrical and optical loss penalties, another prevalent difficulty in slot waveguides. The waveguide core, sandwiched between fingers, is narrowed to 300 nm to localize light in the gap between them by a factor of , as seen in Fig. 2.
Other than the so-called optical field confinement factor (), what matters more is the overlap integral () between optical () and electrical () modal fields. Considering the high , the short effective distance between the two electrodes () brought by the fingers adjacent to the core, the FNS waveguide offers competitive modulation efficiency-loss product (i.e., ) compared to non-slotted waveguide (see Methods). Besides, our simulations revealed that the 10 -long strip-to-finger mode converter does not substantially impact total EOPS loss (conversion efficiency 99.86%). Thanks to its porous structure, FNS topology makes the most of the capillary force action to mitigate material infiltration, another ubiquitous issue in slot waveguides that frequently calls for oxide undercutting [52].
Material handling
The constructed devices must then be aligned with an electric field. Because of its pre-ordered molecular orientation and sensitivity to an external field, FN-LC was found to be fully aligned with > 4 times smaller than required for poling EO polymers [53], in similar structures. The process was carried out at room temperature; no precise temperature monitor was necessary. These relaxed voltage and temperature conditions have additional advantages. First, an inert gas such as N2 is not required around the chip during alignment to avoid oxidation-driven degradation. The FN-LC alignment process is relatively safer because there is no off-gassing. Within the scope of our characterization setup, the device’s performance exhibited little temperature susceptibility thanks to the non-electro-thermal alignment technique. To be more precise, no performance aging was discovered for the FN-LC-coated devices over 150 hours of continuous operation in the range of 10C under full optical and electrical load. EO polymers lose their poling efficiency beyond their glass transition temperature (C). To show that FN-LC is useful at these temperatures, more accelerated aging tests must be done in the future. The lower voltage required for alignment simplifies design concerns for the large-scale integration of FN-LC-enabled photonic devices in which the same RF lines can be utilized.
Characterization
Monitoring optical loss and leakage current throughout the alignment process provides critical information. As seen in Fig. 3(a)-(b), changes in optical loss followed by saturation indicate that domain formation has been completed. In some cases, the change was permanent, indicating that the optic axis continues to remain in a vertical posture. As shown in Fig. 3 (b), the fully-aligned FN-LC MZM showed 2 to 3 dB less material absorption loss compared to the material with unaligned dipoles, which we attribute to the formation of a monodomain over the length of the device and reduced scattering loss (see Methods). The alignment-induced improvement in IL contradicts that seen commonly in EO polymers.
Following the alignment, we performed DC characterization of the optical transmission, as shown in Fig. 3(c), achieving an optical ER of 26 dB and 0.5 V. Various effects, including the birefringence effect, are responsible for this considerable shift [34-35,38,54]. The optical transmission of the proposed MZM shown in Fig. 3(d) reads an IL of 2.1 dB, which we attribute mostly to the propagation loss through the non-optimized subwavelength structure of the deployed FNS waveguides. We also found that FN-LC has approximately the same absorption loss as an available EO polymer at the wavelength of operation.
A more comprehensive figure of merit for a given EOPS is the pure phase efficiency (). The metric is defined as ), where and are the real and imaginary parts of the complex refractive index (), respectively. and loss can be correlated to the estimated values for and . The optical ER, , comprise a comprehensive assessment of EOPS appropriateness. Table LABEL:FOMtable summarizes the abovementioned figure of merits for an EOPS built on hybrid material platforms, including FN-LC and the conventional pn-junction technology. For our proposed FNS waveguide, significantly higher records up to 88.9 have been achieved. We believe the proposed EOPS’ improved phase shift performance stems from combining our device design and the material.
With a high-speed phase shift mechanism, FN-LC might demonstrate significant benefits over traditional PN-LC, in which the slower mechanisms mostly dominate the response. Various aspects of the frequency response characterization of the FN-LC-coated devices are shown in Fig. 4, and the measurement details are covered in Methods. SWG-based grating couplers (GC) were used for light injection into the devices for this step [55-57]. The S-parameter collected by a Vector Network Analyzer (VNA) indicates that 4.18 GHz. This achievement demonstrates a phase shift mechanism with a rise/fall time in the nanosecond range, while PN-LC is limited to functioning in the millisecond range [2]. Accordingly, we employed data and an analytical approach (see Methods) to evaluate the EO properties of the FN-LC material and the proposed EOPS. As a result, an EO coefficient of 24 pmV-1 and an AC modulation efficiency of 25.7 (Vmm) were estimated. The metrics are still significantly lower than the maximum record for EO polymers. This, being the first demonstration, promises further improvement in the future.
Considering the sub-mm phase shifters we used in each arm of the MZM, the mismatch between the optical and RF effective indices is not a limiting factor for the EO bandwidth [18]. The EO bandwidth is primarily limited by the weakly-doped nature of fingers in the fabricated FNS waveguides. As shown by Fig. 1, an RC time constant of GHz is anticipated, which agrees with Fig. 4(d). As an immediate remedy, one could reduce finger lengths (i.e., in Fig. 1) or place electrodes closer to the finger’s end. As a secondary factor, the electrical wire bonds (EWB) that apply the RF signal could also have deteriorated the EO bandwidth. Unlike in slot waveguides, where some optical mode is confined in the two Si rails, less modal field propagates inside fingers in the FNS architecture. A single degree of ion implantation would consequently be sufficient to alleviate the poor shown in this work while maintaining the RF-optical loss trade-off. Besides, as shown in Fig. 2(d), the FNS waveguide’s twin slots reduce the overall . That means, less doping can alternatively achieve a desired , which instead would help improving the device’s (Supplementary Note 3).
We discovered that producing dependable EWBs with less complexity was relatively more straightforward than EO polymer-coated samples. The FN-LC phase shift mechanism appears to swing from a strong birefringence-driven effect at low frequencies to a weaker but faster Pockels effect at higher frequencies. Complex phase manipulation is usually required to develop advanced modulation systems, such as coherent transceivers [58]. Inefficient thermal heaters are frequently used to produce a slow phase shift, as illustrated in Fig. 4(c). Including slow and fast EOPS in a single material could be an appealing feature of the FN-LC-based platform.
Large scale integration
We evaluated the potential large-scale capability of our EOPS structures by creating a SiP chip with 108 MZMs on a 0.5 mm2 area, as depicted in Fig. 5. A subset of devices was photonically wire-bonded (PWB) to an external laser source via a 16-channel fiber array (FA) block for demonstration purposes. The PWB routes were established using adiabatic surface tapers at the chip edge, followed by selective patterning of FN-LC on active components in a post-processing step. Because of its high viscosity, FN-LC can be applied locally, allowing for more controlled integration with the rest of the chip’s components. Thanks to the reduced alignment voltage enabled by utilizing FNS waveguides in our Si-FN-LC design, the step can be paired with a , which is used to bias modulators at their quadrature point. Common EWB and a printed circuit board (PCB) were used to conduct alignment and drive the modulators. The demonstrated FN-LC SiP chip can be viewed as a prototype towards large-scale integration of organic material-enabled PICs, which we credit partly to overcoming the poling difficulties by utilizing the poling-free materials described here. A hypothetical EO polymeric version of this chip necessitates on-chip routing considerations and a separate PCB with potentially higher-voltage handling capabilities for poling.
Towards a truly integrated SiP-FN-LC chip on a large scale, one would require to show the cooperation of several EOPs in relevant scenarios such as the one reported in [59], which require a comprehensive electrical packaging scheme and will be covered in future work. Previously, we have also shown successful integration of other photonic components such as distributed feedback laser [60] and semiconductor laser amplifier [61] using similar photonic packaging technique shown in this work. Furthermore, we anticipate some sealing would be beneficial to accomplishing a properly packaged FN-LC SiP chip despite our early observations of the insensitivity of FN-LC to agents such as humidity and oxygen.
Discussion
We demonstrated GHz-fast EO phase shift and modulation in FN-LC. In contrast to EO polymers, we established that the material does not require any electro-thermal poling processes for alignment. Indeed, the material can be oriented during regular modulator RF operation. An FNS waveguide infiltrated with this material is proposed to reduce the number of lithography steps compared to slot waveguides while benefiting from improved LMI compared to non-slot ones. Using FNS phase shifters as small as 500 in a push-pull MZM, we achieved a DC modulation efficiency of 0.25 Vmm, an optical ER of 26 dB, and an IL of 2.1 dB. The propagation loss of dBmm-1 is still in the order of some of slot waveguides [62-63]. We discovered that the material absorption contributes 0.5 dB of the total IL and is very similar to the case of an EO polymer. The majority of the loss is attributed to the non-optimized FNS waveguide. Reducing fabrication imperfections, including less core sidewall roughness and better infiltration of FN-LC into the two spaces between the core and fingers, can further improve IL. A significant phase shift efficiency in the order of 88.9 was estimated, which we attribute to the material and the FNS waveguide employed in this work. Most importantly, we discovered a high-speed phase shift mechanism with an EO bandwidth > 4.18 GHz that outperforms the stronger but slower birefringence effect found in classic paraelectric-nematic liquid crystals. The EO bandwidth of our FN-LC-covered modulator is limited by the experiment setup and device fabrication, not the suggested architecture or the material, both of which are easy to fix but are out of the scope of this work.
Also, an analytical approach was introduced to estimate the Pockels coefficient of the material, where an AC modulation efficiency of 25.7 Vmm associated with a modest 24 pmV-1 at MHz. As a sufficiently reliable and stable material for practical applications that can scale, the comparatively low modulation efficiency at higher speeds for this first-ever prototype is still a breakthrough and a harbinger of further improvements. Other waveguide embodiments featuring stronger LMI can be coated with FN-LC to achieve high AC modulation efficiencies at the cost of more involved fabrication processes.
The coexistence of a high-speed and effective but slow phase shift mechanism in FN-LC could pave the way for poling-free silicon organic PICs that have similar device metrics to those of EO polymers but require fewer strict steps to be made. The non-thermal nature of the alignment step allows for the selective alignment of phase shifters without affecting adjacent devices, which is one of the main benefits of this yet-to-be-thoroughly validated result. Future research will explore the performance of our proposed poling-free Si-organic modulators based on FN-LC for high-speed data transmission, switching, or transduction at cryogenic temperatures. One hypothesis is to leverage the potentially better material’s elasticity than EO polymers. It could be advantageous to relax the tension caused by different thermal expansions in such a SiP chip without requiring surface passivation or material alterations. Future research will examine the material’s EO properties (i.e., , and ) at cryogenic temperatures.
Methods
Device fabrication: The modulators were patterned on a standard 220 nm silicon on an insulator wafer with a buried oxide layer of 3.5 . An electron beam lithography tool (JBX-8100FS) was used to design the waveguides via positive tone resist (ZEP 520A-7) [56-57]. After that, the devices were dry-etched with an ICP plasma etcher tool (Oxford COBRA). Using maskless lithography (Heidelberg MLA-150), metal electrodes (100 nm gold atop 5 nm titanium as an adhesion layer) were deposited 4.85 m distant from the core waveguide. A passive optical measurement was performed for reference after completing the nanofabrication stages of photonic-only structures and before metallization.
After metallization, both chips were cleaned gently using acetone, IPA, DI water, and a N2 gun, followed by a 30 sec oxygen plasma cleaning for surface passivation just before coating. For the OEO cladding, we used PM-158 (a proprietary FN-LC material optimized for enhanced obtained from Polaris Electro-Optics, Inc.). According to our pre-coating viscosity experiments, the cleaned chips benefitted from being warmed up (70C) to reduce the viscosity of FN-LC during application to the chip. A glass capillary tube was used to apply enough FN-LC to the desired areas. Finally, the chips were cooled to 25∘C before characterization.
FN-LC alignment and DC characterization: While no significant relationship was found between the timing of the alignment stage and the time passed after coating FN-LC, the alignment step was performed on the same day. After coating, the produced samples were mounted on an active EO probe station with temperature control. A source meter unit (Keithley 2400) was employed to apply step voltage and monitor polarization-induced current. This step should be done before AC characterization utilizing the bias Tee shown in Fig. 4(a). As illustrated in Fig. 3(b), we found out a poling field of should be enough for complete alignment, which is at least 4 times smaller than most of existing EO polymers [53,64-65]. The light from a tunable CW laser (Agilent 8164A/81682A) is coupled into grating couplers using an 8-degree polished angle FA. In this phase, we tune the laser wavelength to the highest peak achievable by overlapping the MZM spectrum with that of the GC or PWB. An InGaAs power sensor (Agilent 81635A) detects light. Realtime measurement of optical loss during FN-LC alignment is straightforward and advantageous over EO polymer poling, which requires more advanced control due to elevated temperatures. It can serve as an additional indication of successful alignment. The components of phase shift efficiency (i.e., and ) were extracted using the associated shifts in the vertical () and horizontal () directions as a function of applied voltage.
AC characterization: The same SMU is used as a DC source with a set voltage precision of 0.02 + 24 mV, which is 20 times greater than , thus allowing precise adjustment of MZM bias despite the efficient DC detuning mechanism in FN-LC. An 18 GHz vector network analyzer (Keysigth FieldFox VNA) was added for AC characterization, with port 1 amplified using a microwave amplifier (SHF-s807). It is mixed with a proper amount of DC voltage for optimum MZM performance and simultaneously for FN-LC alignment. A high-speed photodetector (Thorlabs RXM40AF) was employed at the output, along with low-noise amplifiers (picosecond 5828). The VNA data was utilized to evaluate the device S-parameters. The devices were aged at a regulated temperature (25∘C) for 150 hours with no humidity control or hermetic sealing applied, with light ( = 8 dBm) and an RF electrical signal ( = 1.59 V) applied.
The material EO coefficient of FN-LC (i.e., ) was post-processed using the analytical formula [52] (Supplementary Note 1):
where is the calibrated gain of the modulator link, is the total optical loss, is the optical output of the MZM, () is the optical (electrical) input (output) of the detector, is the effective index of the FNS waveguide covered with FN-LC, and is the MZM drive voltage, respectively. Here, and correspond to the detector and the MZM sensitivities. The latter is related to and, within good accuracy, to by:
In Equation 3, denotes the effective distance between the tip of two encountering fingers, which is inversely proportional to the average electric field () over the LMI volume, i.e., where the optical () and electrical () modal fields coexist with the FN-LC. Also, is the overlap integral between and [66] evaluated by:
In Equation 4, n is the field-dependent effective mode index of FNS waveguide, A and V are the cross-section and volume of the LMI region, Z0 is the free space impedance, and is the magnetic field of the optical mode, respectively.
Starting with an approximation (e.g., [67,68]), an electric field simulator is utilized to generate field variable refractive index, which is iteratively called by an optical mode solver (COMSOL Multiphysics). After finite iterations, the solution for the overlap integral should converge to a unique number for each architecture. Both electrical and optical modal field solvers are in 3D to consider the effect of the periodic structure of the FNS waveguide along the light propagation. The semi-double slot waveguide structure of the FNS waveguide can offer compared to of a slot waveguide (=100 nm) (Supplementary Note 2).
However, the characteristic figure of merit of a phase shifter () is . We have also compared a hypothetical nonslotted waveguide with a 300-nm core width surrounded by 2.1 trenches. Our simulations revealed 0.26, 0.27, 0.15 for the proposed FNS waveguide, slotted and nonslotted structures, respectively. If doped properly, one can assume = 200 nm, i.e., 21 times smaller than that of a non-slotted structure. Our 3D point-by-point evaluations of agreed well with this approximation. On the other hand, we have recorded a total propagation loss of dB/mm for the FNS waveguide, versus dB/mm for the above-described, hypothetical nonslotetd structures within the best etching and coating quality we could achieve. Using these three ingredients, one can conclude an improvement of 8.7 times compared to a non-finger-loaded waveguide. The metrics are estimated to be still at least twice higher for the aforementioned slot waveguide thanks to smaller , at the cost of the described challenges.
Data availability
The supporting data for the conclusions of this investigation are available upon reasonable request.
Code availability
The algorithms used for this study are standard and are outlined in “Methods”. Code scripts can be provided by the corresponding authors upon reasonable request.
Acknowledgements
The Natural Sciences and Engineering Research Council of Canada (NSERC), the SiEPICfab consortium, the B.C. Knowledge Development Fund (BCKDF), the Canada Foundation for Innovation (CFI), MITACS, and the Schmidt Sciences funded this research.
Author contribution
This work was conceptualized by I.T., L.C., S.S. C.P., J.S., and N.J. contributed materials and technical information. At the same time, S.J., D.W., and I.T. built the organic Si modulator described in this paper. I.T. conducted measurements while N.J. and O.E. helped with the modulator measurement setup. The manuscript was written by I.T., L.C., and S.S., with assistance from all authors.
Competing interests
The authors declare no competing interests.
e and f in the y-slice perspective (y = 110 nm). Compared to a non-slotted waveguide, the light-matter interactions (LMI)-engineered FNS waveguide acts like a double-slot waveguide to offer 21 times improvement in the as a result of decreased effective electrode distance () as well as 1.75 times higher . () drops to before it overlaps with the encompassing fingers. g A zoomed-in view of indicating that both and are evaluated at each coordinate (x,y,z) in the space, where light interacts with the FN-LC. The marked point proved to be a good reference when estimating (see Methods). h evaluated at 4.18 GHz indicates the RF loss originated by low doping of the FNS waveguide (cm-3).
c A microscopic image of the MZM along with the dipolar order of FN-LC in a section of its finger-loaded non-slotted (FNS) waveguide in the cases of and (other possibilities for dipolar order could exist, since we don’t know what the ground state is with this topology). d Optical output spectrum (normalized) versus applied voltage exhibiting changes in the EOPS complex refractive index ( where for an ideal EOPS). The data is summarized in Table LABEL:FOMtable.
b MZM schematic showing two 500 m-long EOPS and low-loss strip-to-finger-loaded mode converters equipped with GC and PWB optical I/O channels. The alignment has been performed unipolarly, as evidenced by the green arrows. However, as the dark blue arrows show, the RF field is applied bipolarly to achieve push-pull performance. The proposed EOPS made of FN-LC incorporates both slow and fast phase shift mechanisms compared to the conventional approach, shown in c, in which an inefficient thermo-optic effect is required in conjunction with a fast EOPS for advanced modulation schemes. d Measured reveals the existence of a high-speed phase shift mechanism with an EO bandwidth of GHz. Post-processed AC values of based on extracted as a function of frequency are also shown. e A false-colored scanning electron microscope (SEM) image of the Light-matter interaction (LMI)-engineered finger-loaded non-slotted (FNS) waveguide with a zoomed-in view of its mode converter.
[…]
1
Nonlinear optic
2
Paraelectric nematic liquid crystal
3
Ferroelectric nematic liquid crystal