### v3

Abstract
Organic electro-optic materials have demonstrated promising performance in developing electro-optic phase shifters. Their integration with other silicon photonic processes, nanofabrication complexities, and durability remains to be developed. While the required poling step in electro-optic polymers limits their potential and utilization on a large scale, devices made of paraelectric nematic liquid crystals suffer from slow bandwidth. In ferroelectric nematic liquid crystals, we report an additional GHz-fast phase shift that ultimately allows for significant second-order nonlinear optical coefficients related to the Pockels effect. It avoids poling issues and can pave the way for hybrid silicon-organic systems with CMOS-foundry compatibility. We report DC and AC modulation efficiencies of 0.25 Vmm (from liquid crystal orientation) and 25.7 Vmm (from Pockels effect), respectively, an on-chip insertion loss of 2.6 dB, and an electro-optic bandwidth of >4.18 GHz, employing improved light-matter interaction in a waveguide architecture that calls for only one lithography step.
keywords:
Silicon organic hybrid material, electro optic polymer, liquid crystal, phase shifter, modulatorIntroduction
Silicon and silicon nitride photonics have enabled several recent technological advancements, and future roadmaps have been designed based on their capabilities and integration with established CMOS foundries [1]. The electro-optic phase shifter (EOPS) [2] is one of the most fundamental components upon which other essential building blocks in photonic integrated circuits (PIC) are based, including electro-optic (EO) switches [3], EO modulators [4], optical interconnects [5], and comb generators [6]. Traditional phase shift mechanisms often show poor overall performance, such as limited EO bandwidth (BW) and high static power consumption () when leveraging the thermo-optic effect or free-carrier dispersion effect with travelling wave modulators. Free carrier dispersion also leads to doping-induced insertion loss (IL). Innovations in hybrid material platforms promise viable solutions to traditional Si photonics (SiP) shortcomings. However, they typically face significant challenges, such as several complex nanofabrication stages, scaling, and interaction with current CMOS foundries. Modulators based on indium tin oxide (ITO) have recently experienced notable improvements [7]. The additional operations necessary for ITO/insulator thin-film deposition, demanding quality control methods, and their low extinction ratio (ER) suggest further R&D needs [8]. Carrier-free phase shift processes, such as the Pockels effect, are among the more appealing techniques, especially regarding EO bandwidth and energy consumption. These metrics and low IL can excel in LiNbO3 on silicon (LNOI) [9-10]. Due to the low EO coefficient of LiNbO3, however, phase shifters based on LNOI require either a large footprint or a high drive voltage (), where is the voltage necessary to produce shift. A material system based on barium titanate (BTO) can simultaneously meet the requirements of low IL and high EO bandwidth [11-12]. Despite having a significant EO coefficient, BTO is unable to fully utilize the improved light-matter interactions (LMI) provided by some waveguide topologies, such as slot waveguides, failing to give the record-low modulation efficiencies (), where is the EOPS effective length [13]. BTO also involves meticulous fabrication steps, loses its high EO efficiency at cryogenic temperatures, and necessitates process controls for optimum domain development [14].
Furthermore, the preceding approaches cannot meet the trade-offs between various device metrics of a particular modulator, such as footprint, EO and optical BW, IL, , and dynamic energy consumption () [15-16]. Hybrid integration of Si with organic electro-optic (OEO) materials has been proposed to overcome these shortcomings. They have advantages over their inorganic equivalents due to a combination of their EO and physicochemical characteristics. The EO properties of OEO materials, including the Pockels factor () and optical loss, have continuously progressed over the last two decades ( as large as 9000 pmV-1 and absorption coefficients as low as have been reported [17], where is the OEO refractive index and represents the Pockels coefficient). Thanks to their liquid state, an OEO can infiltrate nanostructures during the coating step and benefit from the enhanced LMI. OEO’s low dielectric dispersion [18] promises better modulators and the resultant high modulation efficiencies. As a result, <1 V, sub-mm-long Mach-Zehnder modulators (MZM) with IL < 1 dB, EO bandwidth surpassing 100 GHz, and 100 Jbit-1 have been enabled [19]. Such an all-inclusive, miniaturized modulator can work in extreme temperatures (as high as 383 K (110C) [20] and as low as 4 K [21,22]), which encompass the crucial requirements for a state-of-the-art modulator.
While EO polymers have attracted increased attention mainly because of their remarkable EO characteristics, their application to photonic-integrated circuits (PIC) has challenges. Multiple process steps and often hazardous solvents are needed to prepare EO polymer thin films, necessitating equipment for precise mixing, coating, and drying. In addition, to align chromophores in an EO polymer material, an electro-thermal poling step is necessary. On a larger scale, dedicated DC lines may be required depending on the waveguide geometry and accompanying LMI. Despite considerable breakthroughs in producing thermally and temporally stable polymers, not all poled polymers are resilient to high temperatures ([20]), aging ([23]), and poling-induced optical loss ([24]). For practical applications, the devices may still require a certain level of hermetic sealing to become temperature- and humidity-resistant. EO polymers are also recognized for their lower resistance at higher temperatures, which coincides with detrimental leakage currents. It leads to decreased EO efficiency, which has a direct impact on device performance. Additional nanofabrication steps have been identified as advantageous to resolving this issue, such as exploiting a thin charge-barrier layer of alumina or titania created by atomic layer deposition [25,26]. Poling also limits high-temperature excursions typical of back-end-of-line processing since such excursions would destroy the poling-induced order. The problems above could increase the overall cost, add complexity, and jeopardize the integration of polymer-based hybrid platforms.
Another class of OEO material is PN-LC, which provides a variety of efficient but slow phase shift processes, such as the birefringence effect [27-30]. Ultra-high modulation efficiencies have been reported using strong-LMI implementations, including infiltrated slot waveguides with the material [2]. Unlike polymers, preparing PN-LC devices does not involve a solvent that must be evaporated during post-baking processes. It simplifies the spin coating process and may allow for alternative coating approaches. The material, for example, can be coated on photonic parts via microfluidic channels or fluid dispensing. A PN-LC-based platform is attractive for realizing EO switches, phase shifters, or shutters, but it is not desirable for high-speed applications.
Another state of matter, ferroelectric nematic liquid crystals (FN-LC), has recently been discovered [31-33]. In the landscape of organic EO materials, FN-LC is situated between poled polymers, organic nonlinear optic (NLO) crystals, and PN-LCs, bringing the most desirable aspects of each (see Table 1). Like poled polymers and organic NLO crystals, FN-LC is non-centrosymmetric and has nonzero second-order susceptibilities (). Like organic NLO crystals and PN-LCs, they organize spontaneously–the ordered state is the thermodynamic minimum with significant order parameters. Similar to PN-LCs, the alignment can be controlled over large areas by applying alignment layers at the interface or small external fields. They can also be switched like PN-LCs, with a reorientation of the fast and slow axes in an external field. A unique feature of the FN-LC is that the molecular long axis (slow axis) is highly polarizable and coincident with the director and the polar axis (Fig. 1(a) inset). This creates the condition for achieving very large spontaneous polarizations and large second-order nonlinearities, which have been observed by second harmonic generation measurements [34]. These nonlinearities have been dramatically enhanced by adding NLO dyes to the polar matrix of the FN-LC host material [35]. However, what remains to be demonstrated is a fast EO effect (i.e., Pockels effect) observed in other non-centrosymmetric organic materials. Various operational regimes for different variations of FN-LC were described in [36], where the authors explored other EO properties of the material but did not report the sub-nanosecond characteristic relaxation times or the Pockels effect we demonstrated here. The study found that may be attained with mild external electric fields (). At larger values over 1 V, their model predicted a smaller relaxation time inversely proportional to , down to .
We report demonstrating a considerably faster EO effect in FN-LC, utilizing the nanoscale properties of our SOH structure. We specifically show that FN-LC can be integrated with similar ease as PN-LCs onto a SiP platform, that alignment of the polar axis can be easily achieved over the length of the device, and that a significant EO response can be observed. This paper proposes a hybrid material platform for PICs comprising numerous moderate-speed EOPS that prioritize large-scale scaling, integrability, and ease of fabrication. It includes non-modulation applications, such as switches [3,37], transducers [38], and weight banks [39] for neuromorphic PICs [40] and quantum computing [41].
Results
LMI-engineered design
Waveguides such as Si [42] and plasmonic [43] slotted, sub-wavelength grating (SWG) [44], photonic-crystal [45], and thinned [20], are traditionally employed as EOPS in modulators. However, they often add excessive loss [46], limit optical BW [46], and suffer from poor infiltration and poling [25,47]. Of particular interest for the current work is the extra fabrication stages slotted-based devices demand [38], some of which are not quite CMOS-compatible. For instance, opening oxide cladding over isolated Si islands of SWG or metallic slots of plasmonic waveguides may violate design rules in wafer nanofabrication. Also, exploiting the benefits of structures such as thinned waveguides necessitates taper converters from complete to partial etch Si [20]. Geometries required for efficient and low-loss mode converters, such as strip-to-slot, are not always accessible through nanofabrication runs or have integration challenges with the rest of the PIC. On the other hand, a non-slotted waveguide does not have these issues at the cost of offering a weak LMI [38], thus requiring a high voltage for poling and driving. The device size and EO bandwidth must then be compromised to keep the voltage within the ranges supplied by CMOS drivers.
We propose a finger-loaded strip (FLS) waveguide consisting of two strip-to-finger mode converters to resolve this trade-off, as seen in Fig. 1(a). In analogy to segmented slotted [48] and non-slotted [49] topologies, this structure eliminates the misalignment problem seen in strip-loaded slot waveguides by offering at least one fewer step for lithography and etching. The periodicity and duty cycle of the subwavelength structure of the FLS waveguide is restricted by fabrication capabilities yet selected to produce a better balance between electrical and optical loss penalties, another prevalent difficulty in slot waveguides. The waveguide core, sandwiched between fingers, is narrowed to 300 nm. The optical field confinement factor () of the localized light in the dual gaps between fingertips and the core is shown in Fig. 2(d), and the simulated values are provided in Methods.
Other than the so-called optical field confinement factor , what determines the modulation sensitivity of a device based on the waveguide is the overlap integral () between optical () and electrical () modal fields. Considering the high , the short effective distance between the two electrodes () brought by the fingers adjacent to the core, the FLS waveguide offers competitive modulation efficiency-loss product (i.e., ) compared to a non-slotted waveguide (see Methods). Besides, our simulations revealed that the 10 -long strip-to-finger mode converter does not substantially impact total EOPS loss (conversion efficiency ). Thanks to its porous structure, FLS topology makes the most of the capillary force action to mitigate material infiltration, another ubiquitous issue in slot waveguides that frequently calls for oxide undercutting [50].
Material handling
The constructed devices must then be aligned with an electric field. Because of its pre-ordered molecular orientation and sensitivity to an external field, FN-LC was found to be fully aligned with times smaller than required for poling EO polymers [51], in similar structures. The process was carried out at room temperature; no precise temperature monitor was necessary. These relaxed voltage and temperature conditions have additional advantages. First, an inert gas such as N2 is not required around the chip during alignment to avoid oxidation-driven degradation. The FN-LC alignment process is relatively safer because there is no off-gassing. Within the scope of our characterization setup, the device’s performance exhibited little temperature susceptibility thanks to the non-electro-thermal alignment technique. To be more precise, no performance aging was discovered for the FN-LC-coated devices over 150 hours of continuous operation in the range of 10C under full optical and electrical load. EO polymers lose their poling efficiency beyond their glass transition temperature (C). To show that FN-LC is helpful at these temperatures, more accelerated aging tests must be done in the future. The lower voltage required for alignment simplifies design concerns for the large-scale integration of FN-LC-enabled photonic devices in which the same RF lines can be utilized.
Characterization
Monitoring optical loss and leakage current throughout the alignment process provides critical information. As seen in Fig. 3(a)-(b), changes in optical loss followed by saturation indicate that domain formation has been completed. In some cases, the change was permanent, indicating that the optic axis continues to remain in a vertical posture. As shown in Fig. 3 (b), the fully-aligned FN-LC MZM showed 2 to 3 dB less material absorption loss compared to the material with unaligned dipoles, which we attribute to the formation of a monodomain over the length of the device and reduced scattering loss (see Methods). The alignment-induced improvement in IL is opposite to that seen commonly in EO polymers, where the loss increases with voltage.
Following the alignment, we performed DC characterization of the optical transmission, as shown in Fig. 3(c), achieving an optical ER of 26 dB and 0.5 V. Various processes cause this significant shift, the most prominent of which is the birefringence effect. The optical transmission of the proposed MZM, shown in Fig. 3(d), demonstrates an IL of dB. Based on passive measurements of similar structures coated with a low-loss polymer featuring approximately the same index of refraction, we found dB of the IL primarily attributed to the propagation loss through the non-optimized subwavelength structure of the deployed FLS waveguides and the rest dB stemmed from material absorption at the wavelength of operation. Additionally, the birefringence effect contributes to an increase of around 1 dB, as indicated in Fig.3(d). Nonetheless, just a is required during RF operation, implying an additional loss of dB, resulting in a total IL of dB. We also found that FN-LC has approximately the same absorption loss as an available EO polymer at the wavelength of operation. Aside from fabrication nonidealities, a physical path length difference of integrated into one of the arms has led to a free spectral range (FSR) of , as illustrated by Fig. 3(e). It agrees with the theoretical FSR of nm, where is the group index of the FLS waveguide.
Pure phase efficiency is a seldom reported but very important figure of merit for EOPS for communications applications (). The metric is defined as ), where and are the real and imaginary parts of the complex refractive index (), respectively. and loss can be correlated to the estimated values for and . The optical ER, , comprise a comprehensive assessment of EOPS appropriateness. Table LABEL:FOMtable summarizes the abovementioned figure of merits for an EOPS built on hybrid material platforms, including FN-LC and the conventional pn-junction technology. For our proposed FLS waveguide, significantly higher records up to 88.9 have been achieved, which we credit largely to the material properties.
With a high-speed phase shift mechanism, FN-LC might demonstrate significant benefits over traditional PN-LC, in which the slower mechanisms mostly dominate the response. Various aspects of the frequency response characterization of the FN-LC-coated devices are shown in Fig. 4, and the measurement details are covered in Methods. This step used SWG-based grating couplers (GC) for light injection into the devices [53-55]. The S-parameter collected by a Vector Network Analyzer (VNA) indicates that a -6dB EO bandwidth 4.18 GHz. This achievement demonstrates a phase shift mechanism with a rise/fall time in the nanosecond range, while PN-LC is limited to functioning in the millisecond range [2]. Accordingly, we employed data and an analytical approach (see Methods) to evaluate the EO properties of the FN-LC material and the proposed EOPS. As a result, an EO coefficient of 24 pmV-1 and an AC modulation efficiency of 25.7 (Vmm) were estimated. The metrics are still significantly lower than the maximum record for EO polymers. This, being the first demonstration, promises further improvement in the future.
Considering the sub-mm phase shifters we used in each arm of the MZM, the mismatch between the optical and RF effective indices is not a limiting factor for the EO bandwidth [16]. The EO bandwidth is primarily limited by the weakly doped nature of fingers in the fabricated FLS waveguides. As shown by Fig. 1, a -3dB EO bandwidth () limited by the RC time constant () of GHz is anticipated, which agrees with Fig. 4(d). As an immediate remedy, one could reduce finger lengths (i.e., in Fig. 1) or place electrodes closer to the finger’s end. As a secondary factor, the electrical wire bonds (EWB) that apply the RF signal could also have deteriorated the EO bandwidth. Unlike in slot waveguides, where some optical mode is confined in the two Si rails, less modal field propagates inside fingers in the FLS architecture. A single ion implantation step would consequently be sufficient to alleviate the poor shown in this work while maintaining the RF-optical loss trade-off. Figure 4(d) illustrates superimposed results from applying FN-LC to a properly doped slotted waveguide to validate this hypothesis (data extracted from [67]). Besides, as shown in Fig. 2(d), the FLS waveguide’s twin slots reduce the overall . That means, less doping can alternatively achieve a desired , which instead would help improving the device’s (Supplementary Note 3).
We discovered that producing dependable EWBs with less complexity was relatively more straightforward than EO polymer-coated samples. The FN-LC phase shift mechanism appears to swing from a strong birefringence-driven effect at low frequencies to a weaker but faster Pockels effect at higher frequencies. Complex phase manipulation is usually required to develop advanced modulation systems, such as coherent transceivers [56]. Inefficient thermal heaters are frequently used to produce a slow phase shift (e.g., ), as illustrated in Fig. 4(c). Including slow and fast EOPS in a single material could be an appealing feature of the FN-LC-based platform.
Large scale integration
We evaluated the potential large-scale capability of our EOPS structures by creating a SiP chip with 108 MZMs on a 0.5 mm2 area, as depicted in Fig. 5. A subset of devices was photonically wire-bonded (PWB) to an external laser source via a 16-channel fiber array (FA) block for demonstration purposes. The PWB routes were established using adiabatic surface tapers at the chip edge, followed by selective patterning of FN-LC on active components in a post-processing step. Because of its high viscosity, FN-LC can be applied locally, allowing for more controlled integration with the rest of the chip’s components. Thanks to the reduced alignment voltage enabled by utilizing FLS waveguides in our Si-FN-LC design, the step can be paired with a , which is used to bias modulators at their quadrature point. Common EWB and a printed circuit board (PCB) were used to conduct alignment and drive the modulators. The demonstrated FN-LC SiP chip can be viewed as a prototype towards large-scale integration of organic material-enabled PICs, which we credit partly to overcoming the poling difficulties by utilizing the poling-free materials described here. A hypothetical EO polymeric version of this chip necessitates on-chip routing considerations and a separate PCB with potentially higher-voltage handling capabilities for poling.
The dual-phase shift of FN-LC provides a significant benefit for large-scale integration of dense EOPSs, with an estimated W required to maintain perfect alignment and drive the MZM at the quadrature point. At the frequencies of interest in this work, the proposed MZM’s comparatively short arm length of mm eliminates the necessity for doubly-terminated travelling wave electrodes, another source of static power consumption in MZMs. As a result, the MZM can be driven like a lumped element device with no need for DC routes intended to supply doped heaters. It may also be helpful when numerous devices must coexist in a limited footprint, posing the risk of thermal crosstalk during individual device operations for conventional MZMs. In the case of EO polymers, a scalable method for introducing local heat that does not de-pole neighbouring devices has yet to be established. The poling-free nature of FN-LC mitigates this issue by eliminating the necessity for a locally raised energy level. Each device can be activated using dedicated RF lines routed for normal functioning, which is of special interest for addressing PICs in many applications (e.g., cryogenic), including those with a limited number of total RF and DC I/O lines.
Towards a truly integrated SiP-FN-LC chip on a large scale, one would require to show the cooperation of several EOPs in relevant scenarios such as the one reported in [57], which require a comprehensive electrical packaging scheme and will be covered in future work. Previously, we have also successfully integrated other photonic components, such as distributed feedback laser [58] and semiconductor laser amplifier [59], using a similar photonic packaging technique shown in this work. Furthermore, we anticipate some sealing would be beneficial to accomplishing a properly packaged FN-LC SiP chip despite our early observations of the insensitivity of FN-LC to agents such as humidity and oxygen.
Discussion
We demonstrated GHz-fast EO phase shift and modulation in FN-LC. In contrast to EO polymers, we established that the material does not require any electro-thermal poling processes for alignment. Indeed, the material can be oriented during regular modulator RF operation. An FLS waveguide infiltrated with this material is proposed to reduce the number of lithography steps compared to slot waveguides while benefiting from improved LMI compared to non-slot ones. Using FLS phase shifters as small as 500 in a push-pull MZM, we achieved a DC modulation efficiency of 0.25 Vmm, an optical ER of 26 dB, and an IL of 2.6 dB. The propagation loss of is similar to some slot waveguides [60-61]. We discovered that the material absorption contributes 0.5 dB of the total IL and is very similar to the case of an EO polymer. The majority of the loss is attributed to the non-optimized FLS waveguide. Reducing fabrication imperfections, including less core sidewall roughness and better infiltration of FN-LC into the two spaces between the core and fingers, can further improve IL.
A significant phase shift efficiency in the order of 88.9 was estimated, which we attribute to the material and the FLS waveguide employed in this work. Most importantly, we discovered a high-speed phase shift mechanism of at least 4.18 GHz that outperforms the more substantial but slower birefringence effect in classic paraelectric-nematic liquid crystals. The EO bandwidth of our FN-LC-covered modulator is limited by the experiment setup and device fabrication, not the suggested architecture or the material, both of which are easy to fix but are out of the scope of this work. We established the existence of a Pockels-based linear electro-optic effect in FN-LC materials.
Also, an analytical approach was introduced to estimate the Pockels coefficient of the material, where an AC modulation efficiency of 25.7 Vmm associated with a modest 24 pmV-1 at MHz. The difference between at DC and AC originates from the phase shift mechanism, which is dominant at each regime. At DC, the molecular orientation causes a significant modulation sensitivity (i.e., ). However, the faster Pockels effect dominates at higher frequencies, influenced by the modest reported above. Doping silicon waveguides can also lower , resulting in a higher voltage drop ratio over the LMI region (i.e., , where f is the operating frequency, and represent the resistance of the finger and capacitance of the gap between the core and finger, respectively). The effect is less significant unless at low frequencies, at which , where .
As a sufficiently reliable and stable material for practical applications that can scale, the comparatively low modulation efficiency at higher speeds for this first-ever prototype is still a breakthrough and a harbinger of further improvements. Other waveguide embodiments featuring stronger LMI can be coated with FN-LC to achieve high AC modulation efficiencies at the cost of more involved fabrication processes. The coexistence of a high-speed and effective but slow phase shift mechanism in FN-LC could pave the way for poling-free silicon organic PICs that have similar device metrics to those of EO polymers but require fewer strict steps to be made. The non-thermal nature of the alignment step (i.e., poling-free) allows for the selective alignment of phase shifters without affecting adjacent devices, which is one of the main benefits of this yet-to-be-thoroughly validated result. Future research will explore the performance of our proposed poling-free Si-organic modulators based on FN-LC for high-speed data transmission, switching, or transduction at cryogenic temperatures. The material will likely lose its beneficial DC phase shift as temperature decreases. Besides, alignment should be maintained during the cool-down process. Nonetheless, one may leverage the potentially better material’s elasticity than EO polymers. It could be advantageous to relax the tension caused by different thermal expansions in such a SiP chip without requiring surface passivation or material alterations. Future research will examine the material’s EO properties (i.e., , and ) at cryogenic temperatures.
Methods
Device fabrication: The modulators were patterned on a standard 220 nm silicon on an insulator wafer with a buried oxide layer of 3.5 . The pattern was first corrected for proximity effects using GenIsys Beamer software, applying both short- and long-range proximity effect corrections. The silicon-on-insulator (SOI) sample was cleaned using a standard solvent cleaning process with acetone and isopropyl alcohol (IPA), followed by nitrogen drying and a 30-second oxygen plasma treatment (20 sccm O2 flow). The sample was then baked at 180∘C for dehydration. Zep 520A-7, a positive-tone electron beam lithography (EBL) resist, was spin-coated at 2000 rpm with 1000 rpm/s for 35 seconds to achieve a resist thickness of approximately 260 nm. The sample was subsequently baked at 180∘C to remove any residual solvent. [54-55].
Resist exposure was carried out using a 100 keV EBL system (JEOL JBX-8100FS, JEOL) with a dosage of 213 , a beam current of 1 nA, and a shot pitch of 4 nm. Development was performed in N-Amyl Acetate (ZED-N50, Zeon Specialty Materials Inc.) for 1 minute. The pattern was then transferred into the silicon device layer using an inductively coupled plasma reactive ion etching (ICP-RIE) process (Oxford COBRA). Etching parameters included a chamber pressure of 10 mTorr, a temperature of 15∘C, an ICP power of 600 W, a bias power of 30 W, and a bias voltage of 220 V. The gas mixture consisted of 30 sccm of C4F8 and 25 sccm of SF6, with a total etching duration of 3.5 minutes. Residual Zep resist was removed by exposing the sample to deep ultraviolet light for 15 minutes, followed by immersion in a heated PG remover bath for 15 minutes. The sample was then rinsed in IPA and dried with nitrogen.
For metallization, the sample was treated again with a 30-second oxygen plasma (20 sccm O2 flow) and baked at 110∘C for dehydration. AZ514E photoresist was used in image reversal mode for the lift-off process. Spin coating was performed at 4000 rpm with an acceleration of 1000 rpm/s for 40 seconds. After a post-bake at 110∘C for 1 minute, the sample was exposed using a maskless aligner (MLA-150, Heidelberg Instruments) with a 365 nm wavelength and an exposure dose of 40 mJ/cm2. The sample underwent a second bake at 90∘C for 1 minute, followed by a flood exposure. Development was carried out using a dilute TMAH solution (AZ MIF 300 Developer, MicroChemicals) for 1 minute, followed by a 1-minute rinse in deionized water. Metal electrodes (100 nm gold atop 5 nm titanium as an adhesion layer) were deposited 4.85 m distant from the core waveguide. A passive optical measurement was performed after completing the nanofabrication stages of photonic-only structures and after metallization. No evidence of EO activity was observed before coating FN-LC, which ensures no detectable contribution from free carrier or plasma dispersion effects.
After metallization, the chip is cleaned gently using acetone/IPA and an additional step of DI water for removing any water-soluble residues, dried with N2 gun, and surface-passivated by a step of oxygen plasma treatment for 30 sec just before coating. For the OEO cladding, we used PM-158 (a proprietary FN-LC material optimized for enhanced obtained from Polaris Electro-Optics, Inc.). According to our pre-coating viscosity experiments, the cleaned chip benefitted from being warmed up (70C) to reduce the viscosity of FN-LC during application to the chip. A glass capillary tube was used to apply enough FN-LC to the desired areas. Finally, the chip is cooled to 25∘C before characterization.
AC characterization: The same SMU is used as a DC source with a set voltage precision of 0.02% + 24 mV, which is 20 times greater than , thus allowing precise adjustment of MZM bias despite the efficient DC detuning mechanism in FN-LC. An 18 GHz vector network analyzer (Keysight FieldFox VNA) was added for AC characterization, with port 1 amplified using a microwave amplifier (SHF-s807). It is mixed with a proper amount of DC voltage for optimum MZM performance and simultaneously for FN-LC alignment. A high-speed photodetector (Thorlabs RXM40AF) was employed at the output, along with low-noise amplifiers (picosecond 5828). The VNA data was utilized to evaluate the device S-parameters. The devices were aged at a regulated temperature (25∘C) for 150 hours with no humidity control or hermetic sealing applied, with light ( = 8 dBm) and an RF electrical signal ( = 1.59 V) applied. In Fig. 4(e), an arbitrary function generator drives the amplifier, and the data captured by a digital storage oscilloscope (SDS6204A by SIGLENT) performs fast Fourier transformation.
The material EO coefficient of FN-LC (i.e., ) was post-processed using the analytical formula [52] (Supplementary Note 1)
where is the calibrated gain of the modulator link, is the total optical loss, is the optical output of the MZM, () is the optical (electrical) input (output) of the detector, is the effective index of the FLS waveguide covered with FN-LC, and is the MZM drive voltage, respectively. Here, and correspond to the detector and the modulation sensitivities. The latter is related to and, within good accuracy, to by:
In Equation 3, denotes the effective distance between the tip of two encountering fingers, which is inversely proportional to the average electric field () over the LMI volume, i.e., where the optical () and electrical () modal fields coexist with the FN-LC. Also, is the overlap integral between and [64] evaluated by:
In Equation 4, n is the field-dependent effective mode index of FLS waveguide, A and V are the cross-section and volume of the LMI region, Z0 is the free space impedance, and is the magnetic field of the optical mode, respectively. As a first approximation, one can assume [65,66]. To obtain a more precise value for , an electrical field simulator calculates the local values of , the results of which are called by an optical mode solver (COMSOL Multiphysics) to incorporate the field dependency of . We assumed FN-LC has similar permittivities at the optical and microwave frequencies in analogy to EO polymers [70]. Both electrical and optical modal field solvers are in 3D to consider the effect of the periodic structure of the FLS waveguide along the light propagation. The semi-double slot waveguide structure of the FLS waveguide can offer compared to of a slot waveguide (=100 nm) (Supplementary Note 2).
However, the characteristic figure of merit of a phase shifter () is . We have also compared a hypothetical non-slotted waveguide with a 300-nm core width surrounded by 2.1 trenches. Our simulations revealed 0.26, 0.27, 0.15 for the proposed FLS waveguide, slotted and non-slotted structures, respectively. If doped properly, one can assume = 200 nm, i.e., 21 times smaller than that of a non-slotted structure. Our 3D point-by-point evaluations of agreed well with this approximation. On the other hand, we have recorded a total propagation loss of dB/mm for the FLS waveguide, versus dB/mm for the above-described, hypothetical non-slotted structures within the best etching and coating quality we could achieve. Using these three ingredients, one can conclude an improvement of times compared to a non-finger-loaded waveguide. The metrics are estimated to be still at least twice higher for the aforementioned slot waveguide thanks to smaller , at the cost of the described challenges.
Data availability
The supporting data for the conclusions of this investigation are available upon reasonable request.
Code availability
The algorithms used for this study are standard and are outlined in “Methods.” The corresponding authors can provide code scripts upon reasonable request.
Acknowledgements
The Natural Sciences and Engineering Research Council of Canada (NSERC), the SiEPICfab consortium, the B.C. Knowledge Development Fund (BCKDF), the Canada Foundation for Innovation (CFI), MITACS funding with Dream Photonics Inc., and the Schmidt Sciences funded this research.
Author contribution
This work was conceptualized by I.T., L.C., and S.S. C.P., J.S., J.Y., and N.J. contributed materials and technical information. At the same time, M.H. helped with design and S.J., K.M.A., D.W., and I.T. built the organic Si phase shifters described in this paper. I.T. conducted measurements while N.J. and O.E. helped with the modulator measurement setup. The manuscript was written by I.T., L.C., and S.S., with assistance from all authors.
Competing interests
The authors declare no competing interests.
[…]
1
Nonlinear optic
2
Paraelectric nematic liquid crystal
3
Ferroelectric nematic liquid crystal